OpenStudy (james1769):

Using the completing-the-square method, find the vertex of the function f(x) = 2x2 − 8x + 6 and indicate whether it is a minimum or a maximum and at what point.

OpenStudy (welshfella):

first divide through by 2 = 2( x^2 - 4x + 3) can you continue ?

OpenStudy (james1769):

kinda

OpenStudy (welshfella):

complete the square on what is in the parenthese = 2( x - 2)^2 - A + 3) can you work out what A is?

OpenStudy (james1769):

a is 1?

OpenStudy (welshfella):

the -2 came from dividing the -4 in 4x by 2 Now you have to correct for the last term in the expansion of (x - 2)^2

OpenStudy (welshfella):

(x - 2)^2 = x^2 - 4x + what?

OpenStudy (welshfella):

no, not 1

OpenStudy (welshfella):

(x - 2)(x - 2) = x^2 - 2x - 2x -2 * -2

OpenStudy (welshfella):

sorry guy gotta go right now - the last term is -2*-2 = +4 so A = 4

OpenStudy (james1769):

oh thats right

OpenStudy (james1769):

Maximum at (2, −2) Minimum at (2, −2) Maximum at (2, 6) Minimum at (2, 6) but which one is right?